OpenStudy (anonymous):

5. Solve the following Gas Law Problems. Be sure to include which law applies in the solution. a) A gas has a volume of 100 ml when the pressure is 735 mm. What volume will the gas occupy at 700 mm? b) Change the following temperatures to Kelvin scale: 1) 20 C 2) 85 C 3) - 15 C 4) - 190 C c) Given 90 ml of Hydrogen gas collected with the temperature is 27 degrees C. What volume will the Hydrogen occupy at 42 C? d) A gas measures 500 ml at a temperature of -23 C. Find its volume at 23 C. e) 1.

OpenStudy (anonymous):

I think you'd have more luck with these type of questions in the "Chemistry" group.

OpenStudy (anonymous):

assuming the gas is ideal, we can use the formula: P1*V1 = P2*V2 where P1 = initial pressure V1 = initial volume P2 = final pressure V2 = final volume substituting: 735 * 100 = 700 * (V2) V2 = 105 mL Source: http://www.jiskha.com/display.cgi?id=1301408886

OpenStudy (anonymous):

Same principle formula used on C & D. V1 = initial volume T1 = initial temperature (units in Kelvin) V2 = final volume T2 = final temperature (units in Kelvin) V1*T1 = V2*T2 V1 = 90ml T1 = 300 K V2 = x T2 = 315 K 90*300 = V2 * 315 (90*300) / 315 = x

OpenStudy (anonymous):

thanks!!!! do you think you can help with the other question i posted?

OpenStudy (anonymous):

5 B) http://www.metric-conversions.org/temperature/celsius-to-kelvin.htm D) V1 = 500 T1 = 250 V2 = x T2 = 296 500*250 = x*296 (500*250) / 296 = 422.3mL

OpenStudy (anonymous):

I think A is a test of Boyle's Law whereas C and D are using Charles' Law